OpenStudy (niuuu):

the second term in a geometric sequence is 50. the fourth term in the same sequence is 112.5. what is the common ratio in this sequence?

OpenStudy (kinged):

So, you have the second term, and you multiply that by the common ration to get to the third term then again to get to the forth term, so, what do we get if we divide the fourth term by the second term?

OpenStudy (zzr0ck3r):

\(50,x,112.5\) since it is geometric \(\frac{x}{50}=\frac{112.5}{x}\) Solve for \(x\) and take the positive answer.

OpenStudy (jh99):

x = 75

OpenStudy (jh99):

75 / 50 = ?